OpenStudy (anonymous):

in a geometric sequence, a11=72 and a13=1/2 use the geometric mean to find the value of a12

OpenStudy (anonymous):

\[ \frac{a_{13}}{a_{11}} = \frac{ar^{13-1}}{ar^{11-1}} =r^2 \]

OpenStudy (anonymous):

So \[ r = \sqrt{a_{13}/a_{11}} \]

OpenStudy (anonymous):

\[ a_{13} = ar^{13-1} = arr^{12-1} = ra_{12} \]

OpenStudy (anonymous):

Therefore, \[ a_{12} = \frac{a_{13}}{r} \]

OpenStudy (anonymous):

Well, I mean, you really have two equations here in the end...

OpenStudy (anonymous):

Since \[ a_n=ar^{n-1} \]

OpenStudy (anonymous):

We get: \[ a_{11} = ar^{11-1}\implies 72 = ar^{10}\\ a_{13} = ar^{13-1}\implies \frac12 = ar^{12} \]

OpenStudy (anonymous):

Two equations, two variables.